### v1

# Equality of Hölder exponents for distribution functions of Gibbs measures

Abstract.
We show that the pointwise Hölder spectrum of the distribution function of a Gibbs measure supported on the attractor of a self-conformal subset of the real line satisfies the multifractal formalism.
2020 Mathematics Subject Classification:
37D35, 28A80
1. Introduction
A function $f:\mathbb{R}\to\mathbb{R}$ is said to be Hölder continuous with exponent $\alpha>0$ if there is a constant $C$ such that $|f(x)-f(y)|\leq C|x-y|^{\alpha}$ for all $x$ and $y$. However, this $\alpha$ represents the “worst possible” behavior, and in general, a continuous function can at many points have substantially better regularity than the worst case.
For $\alpha>0$ and $t_{0}\in\mathbb{R}$, write $f\in C^{\alpha}(t_{0})$ if there is a constant $C$ and a polynomial $P$ of degree less than $\alpha$ such that
(1) $|f(t)-P(t-t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t\in\mathbb{R}$}.$
The pointwise Hölder exponent of $f$ at $t$ is the number
(2) $\alpha_{f}(t):=\sup\{\alpha>0:f\in C^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Pointwise Hölder exponents can be difficult to calculate, especially for $\alpha>2$, where, in order to show that $f\not\in C^{\alpha}(t_{0})$, one must prove that no polynomial satisfying (1) exists. For this reason perhaps, many authors use the following, simpler definition of Hölder exponent. Write $f\in\tilde{C}^{\alpha}(t_{0})$ if there is a constant $C$ such that
$|f(t)-f(t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t\in\mathbb{R}$},$
that is, the polynomial $P$ in (1) is constant with value $f(t_{0})$. Define
(3) $\tilde{\alpha}_{f}(t):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Clearly, $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but the reverse inequality may fail in general. For example, if $f(t)=t^{2}$, then $\alpha_{f}(0)=\infty$ (since one can take $P(t)=t^{2}$ in (1) for all $\alpha>2$), but $\tilde{\alpha}_{f}(0)=2$. In addition, a desirable property of Hölder exponents is that they are left unchanged upon perturbation of $f$ by a smooth function $g$. Indeed, $\alpha_{f}(t)=\alpha_{f+g}(t)$, whereas $\tilde{\alpha}_{f}(t)\neq\tilde{\alpha}_{f+g}(t)$ in general.
On the other hand, for certain functions with fractal properties, the two definitions of Hölder exponent coincide. This was shown to be the case by the first author in [1, 2] for certain classes of self-affine functions, including distribution functions of self-similar measures on $\mathbb{R}$ satisfying the open set condition. In this note, we extend this result to distribution functions of arbitrary Gibbs measures on finitely generated self-conformal sets; see the next section for a precise statement.
A function $f$ is multifractal if its pointwise Hölder exponents $\alpha_{f}(t)$ take on more than one value as $t$ ranges through the domain of $f$. For such functions, we are interested in the sets
$E_{f}(\alpha):=\{t\in\mathbb{R}:\alpha_{f}(t)=\alpha\}.$
The Hausdorff and packing multifractal spectra of $f$ are given, respectively, by
$\alpha\mapsto\dim_{H}E_{f}(\alpha)\qquad\mbox{and}\qquad\alpha\mapsto\dim_{P}E_{f}(\alpha),$
where $\dim_{H}$ and $\dim_{P}$ denote Hausdorff and packing dimensions, respectively. If $f=F_{\mu}$ is the distribution function of a Borel probability measure $\mu$, i.e. $F_{\mu}(x):=\mu\left((-\infty,x]\right)$ for $x\in\mathbb{R}$, then by [13, Lemma 5.1] we have
(4) $\tilde{\alpha}_{F_{\mu}}(t_{0})=\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(t_{0},r)\right)}{\log r}.$
For the remainder of this introduction, let $\mu$ a Gibbs measure associated with an arbitrary Hölder potential on the attractor of a conformal iterated function system (IFS) satisfying the open set condition - see the next section for precise definitions. It is well known that the sets
$\mathscr{E}(\alpha;\mu):=\left\{x\in\mathbb{R}\Big{|}\lim_{r\rightarrow 0}\frac{\log\mu\left(B(x,r)\right)}{\log r}=\alpha\right\},\qquad\alpha\in\mathbb{R},$
satisfy the multifractal formalism; that is, $\dim_{H}\mathscr{E}(\alpha;\mu)=\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$, where $\beta^{*}(\alpha)$ is the Legendre transform of a suitable scaling function $\beta(q)$, defined in the next section. The second author and Sumi [13] showed recently that, in the case when all of the maps in the IFS are strictly increasing, the Hausdorff spectrum of the modified level sets
$\mathscr{E}_{*}(\alpha;\mu):=\left\{x\in\mathbb{R}\Big{|}\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(x,r)\right)}{\log r}=\alpha\right\},\qquad\alpha\in\mathbb{R},$
is also given by the multifractal formalism: $\dim_{H}\mathscr{E}_{*}(\alpha;\mu)=\beta^{*}(\alpha)$. It is straightforward to extend this result to general conformal IFS. Thus, as a consequence of our main theorem, the pointwise Hölder spectrum of $F_{\mu}$ is given by
$\dim_{H}E_{F_{\mu}}(\alpha)=\beta^{*}(\alpha).$
In a forthcoming paper, we will show that the situation is different for the packing spectrum.
2. Notation and definitions
Let $X\subset\mathbb{R}$ be a compact interval. A $\mathcal{C}^{1+\epsilon}$ conformal iterated function system on $X$ is given by an index set $I$ with $2\leq\#I<\infty$ and $\mathcal{C}^{1+\epsilon}$ conformal contracting maps $\Phi=(\phi_{i})_{i\in I}$, $\phi_{i}:X\rightarrow X$.
More precisely, we mean that each $\phi_{i}$ extends to a $\mathcal{C}^{1+\epsilon}$ differentiable map on a neighborhood of $X$ which satisfies $0<\left|\phi_{i}^{\prime}(x)\right|<1$ for every $x\in X$ and for each $i\in I$. For ease of presentation, we assume that each of the maps $\phi_{i},i\in I$ is strictly increasing. However, this restriction is not essential. We refer to [18] for further details and basic properties of conformal iterated function systems. Let $\pi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ denote the coding map of $\Phi$ which is for $\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}$ given by
$\bigcap_{n\geq 1}\phi_{\omega_{1}}\circ\dots\circ\phi_{\omega_{n}}(X)=\left\{\pi(\omega)\right\}.$
Let $\Lambda:=\pi(I^{\mathbb{N}})$; then $\Lambda$ satisfies the set equation $\Lambda=\bigcup_{i\in I}\phi_{i}(\Lambda)$. We refer to $\Lambda$ as the self-conformal set generated by $\Phi$. We say that $\Phi$ satisfies the open set condition if $\phi_{i}(\mathrm{Int}(X))\cap\phi_{j}(\mathrm{Int}(X))=\varnothing$ for all $i,j\in I$ with $i\neq j$. To utilize the symbolic thermodynamic formalism (see [6]) we will also need the following definitions. We denote by $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ the left-shift on $I^{\mathbb{N}}$ which becomes a compact metric space endowed with the shift metric. Let $\varphi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be the geometric potential of $\Phi$, given by
(5) $\varphi(\omega):=\log\left(\phi_{\omega_{1}}^{\prime}(\pi(\sigma(\omega)))\right),\quad\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}.$
Since each $\phi_{i}$ is $\mathcal{C}^{1+\epsilon}$ differentiable, it follows that $\varphi$ is Hölder continuous. Let $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be Hölder continuous and denote by
$S_{n}\psi:=\sum_{k=0}^{n-1}\psi\circ\sigma^{k},\qquad n\geq 1$
its ergodic sums. We say that a Borel probability measure $\mu_{\psi}$ on $I^{\mathbb{N}}$ is a Gibbs measure for $\psi$ if there exist two constants $C\geq 1$ and $P\in\mathbb{R}$ such that for each $n\in\mathbb{N}$ and for all $w\in I^{n}$ we have
(6) $C^{-1}\leq\frac{\mu_{\psi}[w_{1},\dots,w_{n}]}{e^{S_{n}\psi(\overline{w})-nP}}\leq C.$
Here $\overline{w}$ denotes the infinite periodic sequence with period block $w$. It is well known that a Gibbs measure $\mu_{\psi}$ for a Hölder continuous potential always exists [6]. Further note that the constant $P$ in the definition of a Gibbs measure is equal to the topological pressure of $\psi$ with respect to $\sigma$, defined by
$\mathcal{P}(\psi):=\lim_{k\to\infty}\frac{1}{k}\log\sum_{w\in I^{k}}\exp S_{k}\psi(\overline{w}).$
We will frequently make use of the fact that any Hölder continuous $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ satisfies the bounded distortion property
(7) $\sup_{n\in\mathbb{N}}\sup_{w\in I^{n}}\sup_{x,y\in I^{\mathbb{N}}}\left|S_{n}\psi(w\rho)-S_{n}\psi(w\tau)\right|<+\infty.$
We remark that if $\mu_{\psi}$ is a Gibbs measure for $\psi$, then it is also a Gibbs measure for $\psi-\mathcal{P}(\psi)$. We may therefore always assume that $\mathcal{P}(\psi)=0$ by replacing $\psi$ with $\psi-\mathcal{P}(\psi)$.
Let
(8) $\alpha_{+}:=\sup_{\omega\in I^{\mathbb{N}}}\limsup_{n\rightarrow\infty}\frac{S_{n}\psi(\omega)}{S_{n}\varphi(\omega)},\qquad\alpha_{-}:=\inf_{\omega\in I^{\mathbb{N}}}\liminf_{n\rightarrow\infty}\frac{S_{n}\psi(\omega)}{S_{n}\varphi(\omega)}.$
It is well known that the multifractal spectrum $\mathcal{F}(\alpha)$ is complete [28], that is, $\mathcal{F}(\alpha)\neq\varnothing$ if and only if $\alpha\in\left[\alpha_{-},\alpha_{+}\right]$. Moreover, the following dichotomy is well known [25]: either we have $\alpha_{-}<\alpha_{+}$, or $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$. Note that in the latter case we have $\alpha_{-}=\alpha_{+}=\dim_{H}\Lambda$.
Since $\varphi<0$, there is for each $q\in\mathbb{R}$ a unique $\beta(q)$ such that
(9) $\mathcal{P}(\beta(q)\varphi+q\psi)=0.$
We denote by
(10) $\beta^{*}(\alpha):=\inf_{q\in\mathbb{R}}\left\{\beta(q)+\alpha q\right\}$
the Legendre transform of $\beta$. If $\alpha_{-}<\alpha_{+}$, then it is well known that $\beta^{*}$ is real-analytic and strictly concave on $(\alpha_{-},\alpha_{+})$ [23].
3. The main theorem
Our main result is the following:
Theorem 3.1.
Let $\Phi=(\phi_{i}:X\rightarrow X)_{i\in I}$ be a conformal iterated function system satisfying the open set condition. Let $\varphi$ denote the geometric potential of $\Phi$ defined in (5) and let $\psi$ be any Hölder potential with $\mathcal{P}(\psi)=0$ such that $\psi$ is not cohomologous to $\varphi$. Let $\mu_{\psi}$ be a Gibbs measure for $\psi$. Then, for the distribution function $F$ of $\mu_{\psi}\circ\pi^{-1}$, we have that $\alpha_{F}=\tilde{\alpha}_{F}$.
Remark 3.2.
The condition that $\psi$ is not cohomologous to $\varphi$ is needed to rule out some degenerate cases. As a simple example, consider the IFS $\Phi=\{x/2,(x+1)/2\}$, whose attractor is the interval $[0,1]$. Clearly, $\varphi\equiv-\log 2$. Choosing $\psi\equiv-\log 2$ as well, we see that $\mu:=\mu_{\psi}\circ\pi^{-1}$ is Lebesgue measure on $[0,1]$, so $F$ is infinitely differentiable on $(0,1)$ and hence $\alpha_{F}(t)\equiv\infty$ on $(0,1)$. On the other hand, it is easy to see that $\tilde{\alpha}_{F}(t)\equiv 1$ on $(0,1)$.
More generally, we claim that for a Hölder continuous $\psi$ with $\mathcal{P}(\psi)=0$ we have that $\psi$ is cohomologous to $\varphi$ if and only if $\Lambda$ is an interval and some/any Gibbs measure $\mu_{\psi}$ for $\psi$ satisfies that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to the Lebesgue measure with a Radon-Nykodym density bounded away from zero and infinity. To prove this, first assume that $\psi$ is cohomologous to $\varphi$. Hence, $P(\varphi)=P(\psi)=0$ and thus, $\dim_{H}(\Lambda)=1$. Hence, by the well-known construction of conformal measures (see e.g. [19, Theorem 4.2.11]), $\Lambda$ has positive Lebesgue measure. Hence, by Lebesgue’s density theorem, we conclude that $X=\bigcup_{i\in I}\Phi_{i}(X)$ and thus, $\Lambda=X$ is an interval. Indeed, if $X\setminus\bigcup_{i\in I}\Phi_{i}(X)$ was non-empty, it would contain an interval of positive length. By standard arguments, this would imply that $\Lambda$ contains no Lebesgue density points, which gives the desired contradiction. Consequently, the normalised Lebesgues measure on $\Lambda$ is a Gibbs measure for $\varphi$. Since $\psi$ is cohomologous to $\varphi$, $\mu_{\psi}$ is also a Gibbs measure for $\varphi$. It follows that $\mu_{\psi}$ is equivalent to the Lebesgue measure with a density bounded away from zero and infinity.
Conversely, assume that $\Lambda$ is an interval and $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to Lebesgue with density bounded away from zero and infinity. Hence, Lebesgue measure is also Gibbs measure for $\psi$. Since Lebesgue measure is also a Gibbs measure for $\varphi$ and $P(\varphi)=P(\psi)=0$, it follows from [19, Theorem 2.2.7] that $\varphi$ is cohomologous to $\psi$.
The key to proving Theorem 3.1 is to prove Proposition 3.4 below. The proof of that proposition uses the following lemma.
Lemma 3.3.
Suppose that $P(\psi)=0$. Suppose that there exists a finite set $A\subset\mathbb{R}$ and $\ell_{0}\geq 1$ such that for all $\ell\geq\ell_{0}$ and all $k_{1},\dots,k_{\ell}\in I$ we have
$\frac{S_{\ell}\psi\left(\overline{k_{1}\dots k_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{k_{1}\dots k_{\ell}}\right)}\in A.$
Then $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $A=\left\{\dim_{H}\Lambda\right\}$.
Proof.
By the bounded distortion property (12), it is easy to see that for each $\ell_{0}\geq 1$,
$\alpha_{-},\alpha_{+}]=\overline{\left\{\frac{S_{\ell}\psi\left(\overline{k_{1}\dots k_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{k_{1}\dots k_{\ell}}\right)}:\ell\geq\ell_{0},k_{1},\dots,k_{\ell}\in I\right\}}.$
Proposition 3.4.
Assume that $\mathcal{P}(\psi)=0$, and $\psi$ is not cohomologous to $\varphi$. Then, for every $k\in\mathbb{N}$ and $x\in\mathbb{R}$ the limit
$\lim_{y\rightarrow x}\frac{F(y)-F(x)}{(y-x)^{k}}$
either does not exist, or it takes the value 0 or $\infty$.
Proof of Proposition 3.4.
The statement is trivial for even $k$, since $F$ is increasing. So below we will assume that $k$ is odd.
Suppose by way of contradiction that
$\lim_{y\rightarrow x}\frac{F(y)-F(x)}{(y-x)^{k}}=L\in(0,+\infty).$
Note that the limit cannot be negative, since $F$ is increasing.
Let $\omega\in I^{\mathbb{N}}$ such that $\pi(\omega)=x$. Let $s_{n},t_{n}]:=\pi\left(\left[\omega_{n}\right]\right)$. Then $s_{n}\leq x\leq t_{n}$, and
(11) $\begin{split}\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}&=\left(\frac{t_{n}-x}{t_{n}-s_{n}}\right)^{k}\cdot\frac{F(t_{n})-F(x)}{(t_{n}-x)^{k}}\\ &\quad+\left(\frac{x-s_{n}}{t_{n}-s_{n}}\right)^{k}\cdot\frac{F(s_{n})-F(x)}{(s_{n}-x)^{k}}.\end{split}$
Denote
$r_{n}:=\frac{t_{n}-x}{t_{n}-s_{n}}.$
Observe that $r_{n}\in[0,1]$ and
$1-r_{n}=\frac{x-s_{n}}{t_{n}-s_{n}}.$
There exists $\ell\in\mathbb{N}$ and $\tau_{1}\dots\tau_{\ell}\in I^{\ell}$ containing at least two different letters such that
$S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)\neq 0.$
Indeed such $\tau_{1}\dots\tau_{\ell}$ always exist for otherwise, Lemma 3.3 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=k=1$, contradicting our hypothesis.
Taking a subsequence if necessary, we may assume that $\lim r_{n}=r\in[0,1]$. Then by (11), we have
$\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}\rightarrow\left(r^{k}+(1-r)^{k}\right)L\in\left[2^{1-k}L,L\right],$
as routine calculus shows that the function $r\mapsto r^{k}+(1-r)^{k}$ is minimized at $r=1/2$. Hence,
(12) $\lim_{n\rightarrow\infty}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)^{k}}=\lim_{n\rightarrow\infty}\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}\in\left[2^{1-k}L,L\right].$
Now, we fix an integer $N\geq 2$ and consider the interval
$\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)=:\left[s_{n}^{N},t_{n}^{N}\right].$
Put
$r_{n}^{N}:=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}.$
Recall that the maps $\phi_{i},i\in I$ are all increasing, so if $u$ and $v$ are two words of the same length and $u\prec v$ in the lexicographical order, then $\pi([u])$ lies to the left of $\pi([v])$.
We claim that there is an infinite subset $\mathcal{N}\subset\mathbb{N}$ such that for each $n\in\mathcal{N}$, the point $x$ and the interval $s_{n}^{N},t_{n}^{N}]$ are separated by at least a full cylinder interval of level at most $n+\ell+1$. To see this, let $j<\ell$ be the smallest integer such that $\tau_{j+1}\neq\tau_{1}$. Such $j$ exists since $\tau_{1}\dots\tau_{\ell}$ contains at least two different letters. Write $\tau_{1}\dots\tau_{\ell}=\tau_{1}^{j}\tau_{j+1}\dots\tau_{\ell}$, and consider two cases:
Case 1. If $\omega$ ends in $\tau_{1}^{\infty}$, then for every sufficiently large $n\in\mathbb{N}$ the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}^{j+1}\tau_{j+1}]\big{)}$ separates $x$ from the interval $\pi\big{(}[\omega|_{n}\tau_{1}\dots\tau_{j+1}]\big{)}$, and hence from $s_{n}^{N},t_{n}^{N}]$.
Case 2. Otherwise, there are infinitely many $n$ such that $\omega_{n+1}\neq\tau_{1}$. Take such an $n$. If $\omega_{n+1}<\tau_{1}$, then
$\omega_{n+1}\dots\omega_{n+\ell+1}\prec\tau_{1}1^{\ell}\prec\tau_{1}\dots\tau_{\ell}\tau_{1},$
so the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}1^{\ell}]\big{)}$ lies between $x$ and $s_{n}^{N},t_{n}^{N}]$. A similar argument, with the inequalities reversed, applies when $\omega_{n+1}>\tau_{1}$.
In either case, the claim follows. By replacing $\mathcal{N}$ by an infinite subset if necessary, we may assume that either $t_{n}^{N}>x$ for all $n\in\mathcal{N}$, or $t_{n}^{N}<x$ for all $n\in\mathcal{N}$. Without loss of generality, we assume the former. The argument in the latter case is similar. Thus, $r_{n}^{N}>0$ for all $n\in\mathcal{N}$.
Let $r_{\min}>0$ be a number such that $|\phi_{i}^{\prime}(x)|\geq r_{\min}$ for all $i\in I$ and $x\in X$. Then
(13) $\mathrm{diam}\pi([uv])\geq\mathrm{diam}\pi([u])\cdot r_{\min}^{|v|}$
for any two finite words $u$ and $v$.
For $n\in\mathcal{N}$, let $u_{n},v_{n}]$ be a cylinder interval of length $n+\ell+1$ separating $x$ and $s_{n}^{N},t_{n}^{N}]$. Then by (13),
$|t_{n}^{N}-x|\geq v_{n}-u_{n}\geq\mathrm{diam}\pi([\omega|_{n}])\cdot r_{\min}^{\ell+1}.$
Hence, there is a constant $C=C(\ell)>1$, depending only on $\ell$ but not on $n$ or $N$, such that
$C^{-1}<\frac{t_{n}^{N}-x}{\mathrm{diam}\big{(}\pi\left(\left[\omega|_{n}\right]\right)\big{)}}<C,$
for all $n\in\mathcal{N}$. We denote this by
(14) $t_{n}^{N}-x\asymp_{\ell}\mathrm{diam}\big{(}\pi\left(\left[\omega|_{n}\right]\right)\big{)}.$
Now (14) implies
(15) $r_{n}^{N}=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}\asymp_{\ell}\frac{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)}{\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}\asymp_{\ell}e^{-NS_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
Hence, by passing to a subsequence, using a diagonal argument, we may assume that $r_{n}^{N}$ converges to some number $r^{N}\in\mathbb{R}$ for every $N\geq 1$.
Using (11) and letting $n\rightarrow\infty$ we obtain
(16) $\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}=\left((r^{N})^{k}+(1-r^{N})^{k}\right)L\geq 2^{1-k}L.$
On the other hand, by the Gibbs property (6) of $\mu_{\psi}$ and the bounded distortion property (12) we have
$\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}$ $=\frac{\mu_{\psi}\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)\right)^{k}}$
$\asymp_{\ell}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}\right]\right)\right)^{k}}\cdot e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}$
$\asymp_{\ell}e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
If $k=1$ this already gives a contradiction, since we then have equality in (16), which implies $S_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=0$.
Suppose now that $k\geq 3$. We have shown that
$2^{1-k}L\leq\big{(}(r^{N})^{k}+(1-r^{N})^{k}\big{)}L=\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}\asymp e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
Since $S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)\neq 0$ it follows that $S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)>0$, and hence, $r^{N}\rightarrow\infty$ as $N\rightarrow\infty$. Since $k$ is odd, $r^{k}+(1-r)^{k}=kr^{k-1}+O(r^{k-2})$ as $r\to\infty$, so we conclude that $(r^{N})^{k}+(1-r^{N})^{k}\asymp(r^{N})^{k-1}$. Thus,
$(r^{N})^{k-1}\asymp e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
Together with (15), this implies
$e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}\asymp(r^{N})^{k-1}\asymp e^{-N(k-1)S_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})},$
and hence,
$S_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=-(k-1)S_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}}),$
i.e.
$S_{\ell}\left(\psi-\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=0.$
We have thus shown that if $\tau_{1}\dots\tau_{\ell}$ contains at least two different letters, then
$S_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})\neq 0\implies S_{\ell}\left(\psi-\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=0.$
Therefore,
(17) $\frac{S_{\ell}\psi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)}\in\left\{1,k\right\}.$
It remains to consider the fixed points $\overline{\tau_{1}}$:
$\frac{S_{\ell}\psi\left(\overline{\tau_{1}}\right)}{S_{\ell}\varphi\left(\overline{\tau_{1}}\right)}=\frac{\psi\left(\overline{\tau_{1}}\right)}{\varphi\left(\overline{\tau_{1}}\right)}.$
Hence, the set
$A:=\left\{\frac{S_{\ell}\psi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)},\quad\ell\geq 1,\quad\tau_{1}\dots\tau_{\ell}\in I\right\}$
contains only finitely many values, at least one of which is 1 or $k$ in view of (17). Now, Lemma 3.3 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=1$. Hence, $\varphi$ is cohomologous to $\psi$, which is a contradiction. ∎
Proof of Theorem 3.1.
We show that, for any $\alpha>1$, if there is a polynomial $P$ of degree $n<\alpha$ such that
(18) $|F(t)-P(t-t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t$},$
then $P$ is constant: $P(t)=f(t_{0})$ for every $t\in\mathbb{R}$. This clearly implies that $\alpha_{F}(t_{0})=\tilde{\alpha}_{F}(t_{0})$.
Suppose, to the contrary, that $P$ has degree $1\leq n<\alpha$ and satisfies (18). We can write
$P(t)=a_{0}+a_{1}t+a_{2}t^{2}+\dots+a_{n}t^{n},$
where $a_{0}=P(0)=F(t_{0})$. We will show by induction that $a_{1}=a_{2}=\dots=a_{n}=0$. First, write
$\frac{F(t)-F(t_{0})}{t-t_{0}}=\frac{F(t)-P(t-t_{0})}{t-t_{0}}+\frac{P(t-t_{0})-F(t_{0})}{t-t_{0}},$
and observe by (18) that
$\left|\frac{F(t)-P(t-t_{0})}{t-t_{0}}\right|\leq C|t-t_{0}|^{\alpha-1}\to 0\qquad\mbox{as $t\to t_{0}$}.$
Hence,
$F^{\prime}(t_{0})=\lim_{t\to t_{0}}\frac{F(t)-F(t_{0})}{t-t_{0}}=\lim_{t\to t_{0}}\frac{P(t-t_{0})-P(0)}{t-t_{0}}=P^{\prime}(0)=a_{1}.$
By Proposition 3.4, it follows that $a_{1}=0$, establishing the basis for the induction.
Next, take $2\leq j\leq n$, and assume we have already shown that $a_{1}=\dots=a_{j-1}=0$. Thus,
$P(t)=F(t_{0})+a_{j}t^{j}+\dots+a_{n}t^{n}.$
Since $j\leq n<\alpha$, we can apply the same idea as above to conclude that
$\lim_{t\to t_{0}}\frac{F(t)-F(t_{0})}{(t-t_{0})^{j}}=\lim_{t\to t_{0}}\frac{P(t-t_{0})-P(0)}{(t-t_{0})^{j}}=\lim_{t\to t_{0}}\sum_{i=j}^{n}\frac{a_{i}(t-t_{0})^{i}}{(t-t_{0})^{j}}=a_{j},$
and therefore, $a_{j}=0$ by Proposition 3.4.
By induction, we have established the claim, so $P(t)\equiv f(t_{0})$, as desired. ∎
4. The pointwise Hölder spectrum of $F_{\mu}$
Recall from the Introduction that $E_{F}(\alpha)=\{t\in\mathbb{R}:\alpha_{F}(t)=\alpha\}$. As a consequence of Theorem 3.1, we obtain the following result.
Corollary 4.1.
With the notation of Theorem 3.1, we have that
$\dim_{H}E_{F}(\alpha)=\beta^{*}(\alpha)\qquad\forall\,\alpha\in[\alpha_{-},\alpha_{+}],$
whereas $E_{F}(\alpha)=\varnothing$ for $\alpha\not\in[\alpha_{-},\alpha_{+}]$.
Proof.
By (4) and Theorem 3.1, it follows that
$E_{F}(\alpha)=\mathscr{E}_{*}(\alpha,\mu_{\psi}).$
As shown in [13], $\dim_{H}\mathscr{E}_{*}(\alpha,\mu_{\psi})=\beta^{*}(\alpha)$ for $\alpha\in[\alpha_{-},\alpha_{+}]$, and $\mathscr{E}_{*}(\alpha,\mu_{\psi})=\varnothing$ otherwise. ∎
Remark 4.2.
The situation is different for the packing spectrum. We illustrate this for the case when $\mu$ is a self-similar measure. Suppose the maps $\{\phi_{i}\}_{i\in I}$ are similarities with contraction ratios $\{r_{i}\}$, and $\mathbf{p}=(p_{1},\dots,p_{m})$ is a probability vector. Put $\psi(\omega)=\log p_{\omega_{1}}$ for $\omega\in I^{\mathbb{N}}$. Likewise we obtain $\varphi(\omega)=\log r_{\omega_{1}}$. Then $\mu_{\psi}$ is just the self-similar measure satisfying
$\mu=\sum_{i\in I}p_{i}\mu\circ\phi_{i}^{-1},$
and the pressure equation (9) becomes
$\sum_{i\in I}r_{i}^{\beta(q)}p_{i}^{q}=1.$
Let $A(x;\mu)$ denote the set of accumulation points of $\log\mu(B(x,r))/\log r$ as $r\to 0$. It was shown by Baek, Olsen and Snigireva [3] under the strong separation condition, and later by Zhou and Chen [30] under the open set condition, that for any closed interval $I$, the set
$\Delta(I):=\{x\in\mathbb{R}:A(x;\mu)=I\}$
has packing dimension
$\dim_{P}\Delta(I)=\max_{\alpha\in I}\beta^{*}(\alpha).$
Let $\alpha_{0}$ be the unique value of $\alpha$ maximizing $\beta^{*}(\alpha)$. Take first $\alpha<\alpha_{0}$, and put $I=[\alpha,\alpha_{0}]$. Since $\mathscr{E}_{*}(\alpha)\supset I$, it follows that
$\dim_{P}E_{F}(\alpha)=\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\geq\dim_{P}\Delta(I)=\beta^{*}(\alpha_{0})=\dim_{P}\Lambda.$
The reverse inequality is trivial since $E_{F}(\alpha)\subset\Lambda$. Hence, $\dim_{P}E_{F}(\alpha)=\beta^{*}(\alpha_{0})$.
For $\alpha\geq\alpha_{0}$, we trivially have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\geq\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$ since $\mathscr{E}_{*}(\alpha;\mu)\supset\mathscr{E}(\alpha,\mu)$. But we also have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\leq\beta^{*}(\alpha)$. This follows directly from a careful examination of the proof that $\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$; see, for instance, [21, p. 501]. We thus conclude that
(19) $\dim_{P}E_{F}(\alpha)=\begin{cases}\beta^{*}(\alpha_{0})&\mbox{if $\alpha_{-}\leq\alpha\leq\alpha_{0}$},\\ \beta^{*}(\alpha)&\mbox{if $\alpha_{0}\leq\alpha\leq\alpha_{+}$}.\end{cases}$
Note that, on the left half of the multifractal spectrum, the Hausdorff dimension does not differentiate between the sets $\mathscr{E}(\alpha;\mu)$ and $\mathscr{E}_{*}(\alpha;\mu)$, but the packing dimension does. Also, it is unusual to see a packing spectrum that is constant on one side of the maximum.
In a forthcoming paper, we will extend the packing spectrum (19) to the case of distribution functions of general Gibbs measures.